The standard heats of formation for CH4, CHCl3, and HCl are -74.8, -132, -92.0 KJ/mole, respectively. Use this information to calculate the heat of reaction for the following:

CH4+3Cl2=CHCl3+3HCl


1
Expert's answer
2021-06-12T10:42:13-0400

"\Delta{H_{reaction}}=\sum{\Delta{H_f({products)}}}-\sum{\Delta{H_f({reactants)}}}=3\times(-92.0)+(-132)-(-74.8)=-333kJ/mol"


Answer: -333 kJ/mol